A club plans to have a fundraiser for which $7 tickets will be sold. The cost of room A rental and refreshments is $200. Find and graph the function p f(n) that gives the 200- profit from the fundraiser when n tickets are sold. Notice that f(O) $200, that is,16 the cost of room rental and refreshments must be paid regardless of how many tickets are sold. How many tickets must be sold to break even (zero profit)? 160 120 What is the function that gives the profit p from the fundraiser when n tickets are sold? 80 (Use integers or decimals for any numbers in the equation.) 12 16 20 24 28 82 36 40 -40 80 120 160 200
Answer #1

For every ticket sold, the Club get $7, meaning after selling n tickets, they havereceived $7n. Since the cost of room and refreshments was $200, this means the profit comes in at p=f(n) = 7n-200. In order to break even:

f(n) = 0 = 7n-200

⇒n=200/7= 28.5714.Since the Club can’t sell, 28.5714 of a ticket, they would need to see 29 functions to break evenrn - 400 - 80 -(20 - 160